# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v2

@@ B.11 Omitted Calculations
$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(y_{t}|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
which yields:
$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (12)
Combining Equations 9, 11, 10 and 12, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
@@
Below we provide two additional results for reference.
Exact Optimization: calculation of the policy performance gap. Since the optimal policy satisfies $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r(y)/\beta)/Z$, we have that:
$V_{r^{\star}}^{\pi^{\star}}=\underset{y\sim\pi^{\star}}{\mathbb{E}}r(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}=\beta\log Z=r^{\star}(y^{\prime})-\beta\log\frac{\pi^{\star}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\;,\;\forall y^{\prime}\in\mathcal{Y}\;.$
Then as motivated by Nika et al. [2024], we can split the performance gap of two-stage RLHF as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
and for DPO we also have:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-r^{\star}(y)+V_{r^{\star}}^{\pi^{\star}}\right]=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
Approximate Optimization: calculation of the reward representation gap. Motivated by Nika et al. [2024] and assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance gap of two-stage RLHF again as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi^{\star}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)+\left(V^{\pi_{\theta}}_{r_{\phi}}-V^{\pi_{\theta}}_{r^{\star}}\right)$ ($\mathcal{F}\subseteq\mathcal{F}_{\Pi}$)
$=\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;,$
and for DPO we have:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $=\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi^{\star}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$\leq\left(V^{\pi^{\star}}_{r^{\star}}-V^{\pi^{\star}}_{\hat{r}_{\theta}}\right)+\left(V^{\pi_{\theta}}_{\hat{r}_{\theta}}-V^{\pi_{\theta}}_{r^{\star}}\right)$
$=\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$

### v3

@@ B.11.5 Formal statement of Theorem 11 and proof
$V_{r^{\star}}^{\pi^{\star}}$ $=\underset{a\sim\pi^{\star},b\sim\pi^{\star}(\cdot|a)}{\mathbb{E}}\left[r^{\star}(a,b)-\beta\log\frac{\pi^{\star}(a,b)}{\pi_{{\mathsf{ref}}}(a,b)}\right]$
$=\beta\log Z$
$=r^{\star}(a^{\prime},b^{\prime})-\beta\log\frac{\pi^{\star}(a^{\prime},b^{\prime})}{\pi_{{\mathsf{ref}}}(a^{\prime},b^{\prime})}\;,\;\forall a^{\prime},b^{\prime}\in\mathcal{V}\;.$
Then we have:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $=\underset{a\sim\pi_{\textup{DPO}},b\sim\pi_{\textup{DPO}}(\cdot|a)}{\mathbb{E}}\left[\beta\log\frac{\pi_{\textup{DPO}}(a,b)}{\pi_{{\mathsf{ref}}}(a,b)}-r^{\star}(a,b)+V_{r^{\star}}^{\pi^{\star}}\right]$
$=\underset{a\sim\pi_{\textup{DPO}},b\sim\pi_{\textup{DPO}}(\cdot|a)}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(a,b)-\beta\log\pi^{\star}(a,b)\right]$
$=\beta\underset{a\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\left(\hat{\theta}_{p,0}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right]+\beta\log\frac{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}\;,$
where $v$ can be any vector in $\mathbb{R}^{d}$. Recall that we require $\pi_{{\mathsf{ref}}}(\cdot|a)$ to be identical for all $a\in\mathcal{V}$ in the proof of Theorem 14. Here we further construct $\pi_{{\mathsf{ref}}}$ to be uniform on the first token. Now observe
$\log\frac{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\underset{a\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}$ $=\log\frac{\sum_{a\in\mathcal{V}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)}{\sum_{a\in\mathcal{V}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)}\;.$
Set $v$ to be $\frac{1}{|\mathcal{V}|}\sum_{a\in\mathcal{V}}\psi(a)$, then we have $\sum_{a\in\mathcal{V}}\left(\psi(a)-v\right)=0$. Since $\Sigma_{\mathcal{D}}$ is already non-singular, we have that $\{\psi(a)-v\}_{a\in\mathcal{V}}$ can span $\mathbb{R}^{d}$. So we can directly apply Lemma 6, and get
$\log\sum_{a\in\mathcal{V}}\exp\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)^{\top}\left(\psi(a)-v\right)\right)-\log\sum_{a\in\mathcal{V}}\exp\left((\hat{\theta}_{p,0})^{\top}\left(\psi(a)-v\right)\right)$
$\geq$ $\langle\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0},\nabla_{\theta}\log\sum_{a\in\mathcal{V}}\exp\left(\theta^{\top}\left(\psi(a)-v\right)\right)|_{\theta=\hat{\theta}_{p,0}}\rangle+\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}$
$=$ $-\underset{a\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\left(\hat{\theta}_{p,0}-(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}})\right)^{\top}\left(\psi(a)-v\right)\right]+\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}\;.$
Therefore, we have
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $\geq\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}$
$=\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}^{2}\|\Sigma_{\mathcal{D}}\|_{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
$\geq\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{2}\left\|\Sigma_{\mathcal{D}}\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right)\right\|_{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
$\geq\frac{\kappa}{2}\langle\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0},\Sigma_{\mathcal{D}}\left(\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right)\rangle\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
$=\frac{\kappa}{2}\left\|\left(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}\right)-\hat{\theta}_{p,0}\right\|_{\Sigma_{\mathcal{D}}}^{2}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}$
$=\Omega\left(\frac{d}{n}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}\right)\;.$
The first inequality comes from Lemma 6; the second equality comes from the non-singularity of $\Sigma_{\mathcal{D}}$; the third inequality comes from a standard property of the spectral norm; the fourth inequality comes from Cauchy-Schwartz inequality; the fifth equality is a simple algebraic equality; and the last equation comes from Theorem 14.
B.12 Omitted Calculations
Calculation of the sub-optimality with respect to the mis-specification error.
First, note that for any $\pi\in\Delta(\mathcal{Y})$, we have:
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi}$ $=\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi^{\star}(y)}{\pi_{{\mathsf{ref}}}(y)}\right]-\underset{y\sim\pi}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]\;,$
$=\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi^{\star}(y)}{\pi(y)}-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]-\underset{y\sim\pi}{\mathbb{E}}\left[r^{\star}(y)-\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}\right]$
$=-\mathsf{KL}\left(\pi^{\star}\|\pi\right)+\underset{y\sim\pi^{\star},y^{\prime}\sim\pi}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
$\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]\;.$
We call it the performance difference lemma (Lemma 1 of Shi et al. [2025]).
For RLHF, we have:
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{RLHF}}}$ $\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi_{\text{RLHF}}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
$\leq\max_{y,y^{\prime}\in\mathcal{Y}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
$\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]}_{\text{reward model mis-specification error}}$
$\quad+\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))-\left(\beta\log\frac{\pi_{\text{RLHF}}(y|x)}{\pi_{{\mathsf{ref}}}(y|x)}-\beta\log\frac{\pi_{\text{RLHF}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]}_{\text{policy model mis-specification error}}\;,$
where the first inequality is by performance difference lemma, and the last two inequalities are by symmetry and the properties of $\max$. And if $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, by the definition of $\pi_{\textup{RLHF}}$, we have
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{RLHF}}}\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]}_{\text{reward model mis-specification error}}\;.$
For DPO, by performance difference lemma, we have:
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\text{DPO}}}$ $\leq\underset{y\sim\pi^{\star},y^{\prime}\sim\pi_{\text{DPO}}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{DPO}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]$
$\leq\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\beta\log\frac{\pi_{\text{DPO}}(y)}{\pi_{{\mathsf{ref}}}(y)}-\beta\log\frac{\pi_{\text{DPO}}(y^{\prime})}{\pi_{{\mathsf{ref}}}(y^{\prime})}\right)\right]}_{\text{policy model mis-specification error}}$
$=\underbrace{\max_{y,y^{\prime}\in\mathcal{Y}}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\text{DPO}}(y)-\hat{r}_{\text{DPO}}(y^{\prime}))\right]}_{\text{surrogate reward model mis-specification error}}\;.$
The first inequality is by performance difference lemma, the second inequality is by symmetry and the property of $\max$, and the last equality is just another interpretation.
Therefore, we can see that the sub-optimality of each algorithm can be upper bounded by the linear model mis-specification error.
@@ Appendix C Implementation Details
More implementation details. For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset. For approximate optimization, we instead compute the empirical BT loss. We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$. During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$. Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.